# AgentNet: Decentralized Evolutionary Coordination for LLM-based Multi-Agent Systems

### v1

@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities in various domains, ranging from basic text understanding to complex reasoning and multimodal integration (OpenAI et al., 2024; Touvron et al., 2023; Yang et al., 2025). Consequently, LLM-based agents have exhibited exceptional performance in numerous tasks, including scientific discovery (Gottweis et al., 2025), automated reasoning (Putta et al., 2024), and website operations (Zhou et al., 2024). However, due to the lack of collective intelligence and collaboration, LLM-based Single Agents struggle to address the complex challenges encountered in the real world. By leveraging collective intelligence through parallel decision-making or workflow collaboration, LLM-based Multi-Agent Systems (MAS) have emerged as a promising framework for tackling complex real-world problems (Guo et al., 2024; Sun et al., 2024; Yang et al., 2024). However, most MAS following the workflow collaboration paradigm rely heavily on a centralized controller or a static, predefined workflow to allocate tasks among agents with fixed roles (Chen et al., 2023a; Hong et al., 2023; Wu et al., 2023; Wang et al., 2024; Ye et al., 2025). While such designs simplify orchestration, they also introduce inherent constraints—including limited scalability, a single point of failure, and challenges to cross-organizational collaboration due to privacy and proprietary knowledge concerns.
@@ 2.1 LLM-based Multi-Agent Systems
The development of LLM-based multi-agent systems (LaMAS) (Yang et al., 2024) has advanced rapidly in recent years. Early frameworks, such as AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2023), made significant strides in establishing foundational architectures for orchestrating multiple LLM agents through structured workflows. AutoGen provided a flexible framework for defining agent interactions, while MetaGPT incorporated software development principles to enhance collaboration. These centralized frameworks proved effective for managing multi-agent interactions. However, they also faced inherent challenges, including limited scalability, single points of failure, and difficulty in dynamically adapting to evolving tasks or incorporating new expertise.
@@ 3.2 Decentralized Network Topology
$w_{m+1}(i,j)=\alpha\cdot w_{m}(i,j)+(1-\alpha)\cdot\sum_{k=1}^{K}S(a_{i}^{m+1},a_{j}^{m+1},t_{m+1}),$ (2)
where $\alpha\in[0,1]$ is a decay factor that balances historical performance with recent interactions, $K$ is the number of recent tasks, and $S(a_{i}^{m+1},a_{j}^{m+1},t_{m+1})$ is a success metric for task $t_{m+1}$ routed from agent $a_{i}^{m+1}$ to $a_{j}^{m+1}$. This adaptive weighting mechanism ensures that the network continuously refines its structure based on operational experience.
@@ 3.4 Dynamic Task Allocation
Each task $t\in T$ is formally represented as a tuple $t=(o_{t},c_{t},p_{t})$, where $d_{t}$ contains the task description in natural language, $c_{t}$ is a vector of capability requirements, and $p_{t}$ denotes the priority level. To efficiently process a new task $t_{m+1}$ after completing task $t_{m}$, AgentNet employs a sophisticated mechanism to select the most suitable initial agent.
@@ 4.1 Experimental Setup
  • Mathematics: This task involves mathematical problem and is evaluated using MATH (Hendrycks et al., 2021a), which includes problems with 7 different types. The training set consists of 100 examples per type (total of 700 problems), while the test set consists of 20 examples per type (total of 140 problems).
  • APPS (Automated Programming Progress Standard): This benchmark assesses more advanced programming problem-solving capabilities, focusing on introductory-level programming problems. The training set contains 400 problems, and the test set includes 100 problems (Hendrycks et al., 2021b).
  • Logical Question Answering: This task tests reasoning and logical question answering abilities using the BBH (Big-Bench Hard) benchmark (Suzgun et al., 2022). The training set follows the MorphAgent setup, selecting 627 examples from 20 tasks. For testing, each task has 5 examples of varying difficulty, totaling 100 test problems.
Metrics
A range of evaluation metrics have been adopted for different tasks. For the Mathematics and the Logical Question Answering tasks, the accuracy metric is utilized to evaluate the consistency of the output answer with the true answer within the specified format. For the Coding task, the average test case pass rate (i.e., the ratio of the number of passed test cases to the total number of test cases) and the ratio of problems passed across all test cases have been employed as the evaluation metrics.
@@
    • Chain of Thought: A prompting technique that elicits step-by-step reasoning from language models (Wei et al., 2023).
    • Synapse: A trajectory-as-exemplar prompting method, which prompts the LLM with complete trajectories of the abstracted states and actions to improve multi-step decision-making. (Zheng et al., 2024)
    • Self-Consistency: A decoding strategy that samples multiple reasoning paths and selects the most consistent answer through majority voting, enhancing reliability (Wang et al., 2023).
    • Self-Refinement: An iterative approach where models critically evaluate and improve their own solutions over multiple passes, progressively enhancing solution quality (Madaan et al., 2023).
@@
    • MetaGPT: A software development framework where specialized agents (like product manager, architect, engineer) collaborate in a waterfall workflow to complete complex engineering tasks (Hong et al., 2023).
    • AFLOW: A framework that optimizes agent workflows using Monte Carlo Tree Search over code-represented workflows with execution feedback (Zhang et al., 2024c).
    • MorphAgent: A framework featuring self- evolving agent profiles that dynamically optimize individual expertise in the profile through three metrics (Lu et al., 2024).
@@ 4.2 Main Results
Table 1: Performance Comparison of Different Methods Across Various Tasks. The evaluation metrics are accuracy for Math and Logical QAs tasks, and (average test case pass rate, problem pass ratio) for the Coding tasks.
Backbone Category Method MATH APPS BBH
GPT-4o-mini Single Agent Direct 31.43 (21.43, 20.00) 59.00
React 55.71 (21.87, 17.00) 80.00
Synapse 77.14 (75.88, 60.00) 79.00
Self-Consistency 54.28 (52.41, 43.00) 85.00
Self-Refinement 68.57 (44.90, 34.00) 81.00
Multi-Agent MegaGPT 73.57 (70.93, 56.00) 53.00
AFLOW 85.00 (49.36, 37.00) 75.00
MorghAgent 80.71 (0, 0) 56.00
AgentNet 85.00 (70.59, 58.00) 86.00
Table 1 presents a performance comparison across various tasks: Math, Coding, and Logical QAs. The evaluation metrics include accuracy for Math and Logical QAs tasks, and the average test case pass rate and problem pass ratio for Coding tasks. When compared to single-agent methods, AgentNet performs at par or better across all tasks. While Synapse and React achieve notable results in Math and Logical QAs, AgentNet’s decentralized, multi-agent setup provides superior flexibility and performance, especially in more complex tasks. In comparison to other multi-agent systems, AgentNet demonstrates clear advantages. MegaGPT, with its centralized control, struggles with scalability and flexibility, showing weaker performance in tasks like Logical QAs (53.00 accuracy) and Coding (APPS). Notably, MorghAgent struggles with understanding the problem in APPS tasks, as it generates its own dataset and test cases during training. This self-generated data causes morphagent to produce results that do not compile properly.
@@ 4.3 Ablation Study
Based on the observed results in Table  2, the experiments demonstrate a clear performance improvement when using AgentNet compared to the baseline without warm-up. In the MATH task, AgentNet outperforms the non-warm-up method by a significant margin, achieving a score of 85.00 compared to 77.86. Similarly, in the APPS task, AgentNet shows an increase in both accuracy and performance score, from (0.5911, 0.44) to (0.6782, 0.54). Finally, in the BBH task, the accuracy improves from 76% to 86%, confirming that AgentNet’s ability to autonomously evolve through experience leads to more optimized performance in various tasks.
Table 2: Performance Comparison of AgentNet vs. Without Warm-up
3 Agents MATH APPS BBH
w/o warm-up 77.86 (59.11, 44.00) 76.00
AgentNet 85.00 (67.82, 54.00) 86.00
These results highlight the crucial role of adaptive learning in AgentNet, which enables agents to refine their capabilities and improve performance over time. The substantial improvement in all tasks suggests that the warm-up phase, through its adaptive learning process, allows agents to better specialize and optimize their performance, confirming the importance of this mechanism for the system’s overall efficiency and task handling.
@@ 5.1 Scalability and Robustness of the System
Based on the experimental results as illustrated in Figure  7, we observed that both training and testing performance improve slightly as the number of agents and executor pool limit increase. However, the improvements are incremental, with diminishing returns as the system scales up. Specifically, performance in the training phase increased from 80.38 for 3 agents and 30 executors to 81.18 for 9 agents and 40 executors. In the testing phase, performance fluctuated between 80 and 86, with the highest performance seen in configurations with 40 executors.
@@ 5.2 Autonomous Specialization of Agents
Based on the observed results in Figure  8, the experiments demonstrate that AgentNet’s multi-agent system can naturally specialize agents in a decentralized environment. With varying numbers of agents and a fixed executor pool of 40 pieces, the ability scores across different tasks such as reasoning, language, knowledge, and sequence showed significant variation. As the number of agents increased, specialization became more evident, particularly in complex tasks, with certain agents excelling in specific areas while others focused on different abilities. This highlights AgentNet’s capacity to dynamically refine agent expertise and optimize performance in a decentralized, task-driven system.
@@
5.3 Limitations and Future Work
Despite AgentNet implementing a fully distributed, adaptive learning multi-agent system (MAS) with dynamic task allocation, several important limitations remain that require further exploration in future work. One key challenge is how to improve task performance in heterogeneous agent environments. In real-world applications, agents often vary significantly in terms of model capabilities, workflow structures, tools, and available data. The impact of such heterogeneity on AgentNet’s performance, especially in terms of task coordination and resource allocation, remains an open question. Understanding how to adapt the system to handle such variations efficiently will be crucial for its scalability and effectiveness in complex environments. Secondly, the decision-making process of the router within each agent, particularly in relation to exploration and discovery, requires more in-depth study. Currently, the router selects agents from a relatively small pool of predefined candidates. However, in larger-scale systems involving hundreds or potentially thousands of agents, the challenge of accurately identifying the most suitable agent for task delegation becomes significantly more complex. This problem is further compounded in heterogeneous settings, where different agents may possess distinct strengths and weaknesses. To address this, future research could focus on developing more sophisticated routing mechanisms that can autonomously identify and delegate tasks to the most appropriate agents, even in large and diverse agent pools. Additionally, a promising direction for future work involves designing incentives that encourage the router to explore agents beyond the predefined candidate set. By enabling AgentNet to dynamically discover new agents or specialized capabilities, such an approach would enhance its adaptability and scalability, ultimately improving the system’s overall performance and autonomy.

### v2

@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable capabilities in various domains, ranging from basic text understanding to complex reasoning and multimodal integration (OpenAI et al., 2024; Touvron et al., 2023; Yang et al., 2025). Consequently, LLM-based agents have exhibited exceptional performance in numerous tasks, including scientific discovery (Gottweis et al., 2025), automated reasoning (Putta et al., 2024), and website operations (Zhou et al., 2024). However, due to the lack of collective intelligence and collaboration, LLM-based Single Agents struggle to address the complex challenges encountered in the real world. By leveraging collective intelligence through parallel decision-making or workflow collaboration, LLM-based Multi-Agent Systems (MAS) have emerged as a promising framework for tackling complex real-world problems (Guo et al., 2024; Sun et al., 2024; Yang et al., 2024). However, most MAS following the workflow collaboration paradigm rely heavily on a centralized controller or a static, predefined workflow to allocate tasks among agents with fixed roles (Chen et al., 2023a; Hong et al., 2024; Wu et al., 2023; Wang et al., 2024; Ye et al., 2025). While such designs simplify orchestration, they also introduce inherent constraints—including limited scalability, a single point of failure, and challenges to cross-organizational collaboration due to privacy and proprietary knowledge concerns.
@@ 2.1 LLM-based Multi-Agent Systems
The development of LLM-based multi-agent systems (LaMAS) (Yang et al., 2024) has advanced rapidly in recent years. Early frameworks, such as AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2024), made significant strides in establishing foundational architectures for orchestrating multiple LLM agents through structured workflows. AutoGen provided a flexible framework for defining agent interactions, while MetaGPT incorporated software development principles to enhance collaboration. These centralized frameworks proved effective for managing multi-agent interactions. However, they also faced inherent challenges, including limited scalability, single points of failure, and difficulty in dynamically adapting to evolving tasks or incorporating new expertise.
@@ 3.2 Decentralized Network Topology
$w_{m+1}(i,j)=\alpha\cdot w_{m}(i,j)+(1-\alpha)\cdot S(a_{i}^{m+1},a_{j}^{m+1},t_{m+1}),$ (2)
where $\alpha\in[0,1]$ is a decay factor that balances historical performance with recent interactions, and $S(a_{i}^{m+1},a_{j}^{m+1},t_{m+1})$ is a success metric for task $t_{m+1}$ routed from agent $a_{i}^{m+1}$ to $a_{j}^{m+1}$. This adaptive weighting mechanism ensures that the network continuously refines its structure based on operational experience.
@@ 3.4 Dynamic Task Allocation
Each task $t\in T$ is formally represented as a tuple $t=(o_{t},c_{t},p_{t})$, where $o_{t}$ contains the task description in natural language, $c_{t}$ is a vector of capability requirements, and $p_{t}$ denotes the priority level. To efficiently process a new task $t_{m+1}$ after completing task $t_{m}$, AgentNet employs a sophisticated mechanism to select the most suitable initial agent.
@@ 4.1 Experimental Setup
  • Mathematics: This task involves mathematical problem and is evaluated using MATH Hendrycks et al. (2021), which includes problems with 7 different types. The training set consists of 100 examples per type (total of 700 problems), while the test set consists of 20 examples per type (total of 140 problems).
  • Logical Question Answering: This task tests reasoning and logical question answering abilities using the BBH (Big-Bench Hard) benchmark Suzgun et al. (2022). The training set follows the MorphAgent setup, selecting 627 examples from 20 tasks. For testing, each task has 5 examples of varying difficulty, totaling 100 test problems.
  • Function-Calling: This benchmark evaluates the agent’s ability to perform tool-augmented task planning and API usage, based on the API-Bank dataset Li et al. (2023). We construct a training set of 100 tasks and a test set of 100 tasks, randomly sampled from the full API-Bank corpus. Since the original dataset does not include category labels, we annotate each task using GPT-4o-mini to assign one of the seven task types: health, account, schedule, information, housework, finance, and others. Each task is further categorized into one of three difficulty levels, determined by prompt complexity and required toolchain length.
@@
    • React: A prompting technique that elicits step-by-step reasoning from language models Wei et al. (2023); Yao et al. (2023).
    • Synapse: A trajectory-as-exemplar prompting method, which prompts the LLM with complete trajectories of the abstracted states and actions to improve multi-step decision-making. Zheng et al. (2024)
    • Self-Consistency: A decoding strategy that samples multiple reasoning paths and selects the most consistent answer through majority voting, enhancing reliability Wang et al. (2023).
    • Self-Refinement: An iterative approach where models critically evaluate and improve their own solutions over multiple passes, progressively enhancing solution quality Madaan et al. (2023).
@@
    • MorphAgent: A framework featuring self- evolving agent profiles that dynamically optimize individual expertise in the profile through three metrics Lu et al. (2024).
    • MetaGPT: A software development framework where specialized agents (like product manager, architect, engineer) collaborate in a waterfall workflow to complete complex engineering tasks Hong et al. (2024).
    • AFLOW: A framework that optimizes agent workflows using Monte Carlo Tree Search over code-represented workflows with execution feedback Zhang et al. (2024c).
    • GPTSwarm: A framework modeling agents as computational graphs with automatic optimization of both prompts and agent collaboration patterns Zhuge et al. (2024).
@@ 4.2 Main Results
Table 1: Performance comparison of different methods across various tasks. In all multi-agent methods, we set 3 agents for each method to ensure a fair comparison. The best result is in bold, while the second is underlined.
Backbone Category Method MATH(Acc/%) BBH(Acc/%) API-Bank(Acc/%)
DeepSeek-V3 Single Agent Direct 47.86 69.00 26.00
React 77.14 88.00 29.00
Synapse 89.28 92.00 28.00
Self-Consistency 88.00 85.00 29.00
Self-Refinement 87.14 84.00 25.00
Multi-Agent MorghAgent 39.29 56.00 16.00
MetaGPT 92.14 64.00 22.00
AFLOW 91.67 88.00 28.00
GPTSwarm 72.14 90.00 21.00
AgentNet 92.86 94.00 30.00
GPT-4o-mini Single Agent Direct 31.43 59.00 15.00
React 55.71 80.00 24.00
Synapse 77.14 79.00 22.00
Self-Consistency 54.28 85.00 22.00
Self-Refinement 68.57 81.00 23.00
Multi-Agent MorghAgent 80.71 56.00 16.00
MetaGPT 73.57 53.00 19.00
AFLOW 85.00 75.00 21.00
GPTSwarm 85.00 86.00 13.00
AgentNet 85.00 86.00 29.00
Qwen-turbo Single Agent Direct 37.85 57.00 27.00
React 53.57 69.00 23.00
Synapse 67.14 68.00 24.00
Self-Consistency 64.28 70.00 28.00
Self-Refinement 76.43 74.00 23.00
Multi-Agent MorghAgent 16.43 56.00 9.00
MetaGPT 63.57 51.00 20.00
AFLOW 82.14 57.00 22.00
GPTSwarm 79.29 75.00 30.00
AgentNet 81.43 92.00 32.00
Table 1 summarizes performance across Math, Logical QA tasks and API-Calling tasks. For Math, Logical QAs and API-Calling tasks, accuracy is reported. Compared to single-agent methods (e.g., Synapse, ReAct), AgentNet achieves competitive or superior performance across all tasks. While ReAct performs well on Math and Logic, its static prompting strategy limits generalization to more complex tasks. Against multi-agent baselines, AgentNet consistently outperforms centralized frameworks such as MetaGPT, which suffers from limited scalability—e.g., only 53.00% accuracy on Logical QA. AgentNet’s decentralized coordination and retrieval-augmented memory contribute to its robustness across domains, particularly in tasks requiring contextual understanding and adaptive role specialization.
4.3 Experiments on Heterogeneous Agents
To investigate the impact of agent diversity on performance, we designed a heterogeneity experiment across different settings on the BBH task. Agents were tested under four configurations: fully homogeneous (identical models and capabilities), LLM heterogeneity (different language models, same capabilities), skill heterogeneity (same model but varied capabilities), and a combination of both. This design allows us to isolate and analyze how model-level and capability-level diversity influence multi-agent collaboration.
Table 2: BBH Accuracy under Different Heterogeneity Settings (Acc%)
Setting Fully Homogeneous Skill Hetero. LLM Hetero. Both Hetero.
3 Agents 0.86 0.84 0.81 0.81
5 Agents 0.79 0.86 0.85 0.85
The results show that the impact of heterogeneity on multi-agent performance depends on team size. With 3 agents, the fully homogeneous setting performs best, while introducing either model or skill diversity reduces accuracy, suggesting uniform reasoning is more effective in small teams. However, with 5 agents, heterogeneous configurations outperform the homogeneous one, indicating that diversity enhances collaboration and complementary reasoning in larger teams. Overall, heterogeneity may introduce coordination overhead in small groups but offers clear benefits at larger scales.
@@ 4.4 Ablation Study
Router Effectiveness in AgentNet
To evaluate AgentNet’s decentralized router, experiments were conducted comparing AgentNet with ablation configurations: "Totally Random", "Random Operations", "Random Next Agent ID" and a centralized "Global Router". Each router manages external routing (selecting the next agent) and internal routing (deciding to forward, split, or execute).
Performance was tested on the BBH task (training: 627 problems, testing: 100 problems), with results in Figure 5. AgentNet outperforms randomized methods, achieving 82.14% accuracy during training and 86.00% during testing. Randomizing operations (forward/split/execute) affects task execution more directly, randomizing next agent ID primarily results in suboptimal task delegation but does not disrupt task completion as severely. These results underscore the critical role of effective routing and suggest that optimizing routing decisions can significantly enhance multi-agent system performance.
Figure 5: AgentNet’s Router Performance on the BBH (Backbone: gpt-4o-mini)
@@
Results in Table 3 clearly indicate that AgentNet significantly improves performance compared to the non-evolution baseline. On the MATH task, AgentNet achieves a score of 85.00 versus 77.86. For the function-calling task, performance improves notably from 23.00 to 32.00. On the BBH task, accuracy rises from 76% to 86%, demonstrating the impact of evolution phase for AgentNet.
Table 3: Performance Comparison of AgentNet vs. Without evolution (Backbone: gpt-4o-mini)
3 Agents MATH API-Bank BBH
Acc(%) Acc(%) Acc(%)
w/o evolution 77.86 23.00 76.00
AgentNet 85.00 32.00 86.00
These results confirm that AgentNet’s adaptive learning during the evolution phase effectively enhances agent specialization and task performance, demonstrating its essential role in the system’s optimization and overall efficiency.
@@ 5.1 Scalability and Robustness of the System
Based on the experimental results as illustrated in Figure  8, we observed that both training and testing performance improve slightly as the number of agents and executor pool limit increase. However, the improvements are incremental, with diminishing returns as the system scales up. Specifically, performance in the training phase increased from 80.38 for 3 agents and 30 executors to 81.18 for 9 agents and 40 executors. In the testing phase, performance fluctuated between 80 and 86, with the highest performance seen in configurations with 40 executors.
@@ 5.2 Autonomous Specialization of Agents
Based on the observed results in Figure  9, the experiments demonstrate that AgentNet’s multi-agent system can naturally specialize agents in a decentralized environment. With varying numbers of agents and a fixed executor pool of 40 pieces, the ability scores across different tasks such as reasoning, language, knowledge, and sequence showed significant variation. As the number of agents increased, specialization became more evident, particularly in complex tasks, with certain agents excelling in specific areas while others focused on different abilities. This highlights AgentNet’s capacity to dynamically refine agent expertise and optimize performance in a decentralized, task-driven system.
@@
5.3 Evolution of Agents Networks
The evolution of the agent network in our experiment is illustrated in Figure 10, which demonstrates the transition of a multi-agent system composed of 5 agents running on the BBH (627 pieces) benchmark. The figure captures the network at three key stages: the initial state, an intermediate state, and the final evolved state.
Figure 10: Evolution Example of Agents Networks.
In the initial state, the network is fully connected with uniform connection values of 1.00, indicating equal capabilities among all agents. At this stage, there is no specialization, and all agents are equally equipped to handle tasks.
As the network evolves, agents begin to specialize, and connection values vary, reflecting the strength of collaboration. Stronger connections indicate tighter cooperation, while weaker ones suggest less interaction. This evolution shows how agents naturally adapt and form more efficient collaboration patterns.
By the final stage, the network exhibits clear specialization, with agents taking on distinct roles. The connection values further emphasize the growing cooperation between specialized agents, improving task performance. This progression demonstrates the effectiveness of decentralized coordination, where evolving collaboration enhances task allocation, scalability, and fault tolerance.
6 Limitations and Future Work
Despite AgentNet implementing a fully distributed, adaptive learning multi-agent system (MAS) with dynamic task allocation, several important limitations remain that require further exploration in future work.
One key challenge is how to improve task performance in heterogeneous agent environments. In real-world applications, agents often vary significantly in terms of model capabilities, workflow structures, tools, and available data. The impact of such heterogeneity on AgentNet’s performance, especially in terms of task coordination and resource allocation, remains an open question. Understanding how to adapt the system to handle such variations efficiently will be crucial for its scalability and effectiveness in complex environments.
Secondly, the decision-making process of the router within each agent, particularly in relation to exploration and discovery, requires more in-depth study. Currently, the router selects agents from a relatively small pool of predefined candidates. However, in larger-scale systems involving hundreds or potentially thousands of agents, the challenge of accurately identifying the most suitable agent for task delegation becomes significantly more complex. This problem is further compounded in heterogeneous settings, where different agents may possess distinct strengths and weaknesses. To address this, future research could focus on developing more sophisticated routing mechanisms that can autonomously identify and delegate tasks to the most appropriate agents, even in large and diverse agent pools.
Additionally, a promising direction for future work involves designing incentives that encourage the router to explore agents beyond the predefined candidate set. By enabling AgentNet to dynamically discover new agents or specialized capabilities, such an approach would enhance its adaptability and scalability, ultimately improving the system’s overall performance and autonomy.
@@ 7 Conclusion
Acknowledgement
The authors gratefully acknowledge the generous support provided by Bytedance, which has been instrumental in enabling the successful execution of this research. Their sponsorship significantly contributed to the advancement and completion of this work.
Acknowledgement
This research was supported by Bytedance through a sponsored project that facilitated the execution and completion of this work.
Ethics Statement
This study did not involve human participants, animal subjects, or the use of personal data. All datasets and benchmarks employed are publicly available and used in accordance with their respective licenses. Therefore, no ethics approval was required.